# A Principle of Targeted Intervention for Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Steering cooperative multi-agent reinforcement learning (MARL) towards desired outcomes is challenging, particularly when the global guidance from a human on the whole multi-agent system is impractical in a large-scale MARL. On the other hand, designing mechanisms to coordinate agents most relies on empirical studies, lacking a easy-to-use research tool. In this work, we employ multi-agent influence diagrams (MAIDs) as a graphical framework to address the above issues. First, we introduce interaction paradigms that leverage MAIDs to analyze and visualize existing approaches in MARL. Then, we design a new interaction paradigm based on MAIDs, referred to as targeted intervention that is applied to only a single targeted agent, so the problem of global guidance can be mitigated. In our implementation, we introduce a causal inference technique—referred to as Pre-Strategy Intervention (PSI)—to realize the targeted intervention paradigm. Since MAIDs can be regarded as a special class of causal diagrams, a composite desired outcome that integrates the primary task goal and an additional desired outcome can be achieved by maximizing the corresponding causal effect through the PSI. Moreover, the bundled relevance graph analysis of MAIDs provides a tool to identify whether an MARL learning paradigm is workable under the design of an interaction paradigm. In experiments, we demonstrate the effectiveness of our proposed targeted intervention, and verify the result of relevance graph analysis.
@@ 1 Introduction
To address these MARL challenges, researchers often employ guidance under various interaction paradigms to enhance performance, efficiency and applicability. Intrinsic rewards, for instance, can improve sample efficiency and direct agent behaviour by encouraging exploration or coordination jaques2019social ; du2019liir ; mohamed2015variational . Another approach incorporating human feedback, leverages human expertise to accelerate learning. This helps agents tackle objectives that are complex to specify via simple reward functions, and aligns agent behaviour to desired norms or human values christiano2017deep ; hu2022human ; mnih2015human .
All these interaction paradigms cannot be simply categorized by the conventional MARL learning paradigms (i.e., independent learning and centralized training).
@@
The main contributions of this paper are summarized as follows: (1) We formalize the MARL learning process in the language of MAIDs. Based on MAIDs, we categorize the existing MARL interaction paradigms, and then propose a new one referred to as the targeted intervention paradigm. In this new interaction paradigm, only a single targeted agent is intervened on. (2) The solvability of existing MARL learning paradigms under various MARL interaction paradigms can be analyzed, with the help of visualizing the corresponding relevance graphs. (3) Since MAIDs can be viewed as causal diagrams, we draw on the principle of causal inference to implement the targeted intervention paradigm, which we refer to as the Pre-Strategy Intervention (PSI). It is provable that the PSI can reach a composite desired outcome consisting of the primary task goal and an additional desired outcome. (4) The PSI is implemented as a pre-policy module that can be integrated into generic MARL algorithms.
@@ 3 Multi-Agent Reinforcement Learning Interaction Paradigms
This paper focuses on analyzing interaction paradigms in MARL with guidance and proposes a novel approach for the targeted intervention paradigm. For a comprehensive understanding, we first formalize the two existing interaction paradigms: direct interaction and global intervention strouse2022collaborating ; mosqueira2023human ; wang2025m3hf , using multi-agent influence diagrams (MAIDs). Subsequently, we will elaborate on the details of our proposed targeted intervention paradigm.
@@
Figure 1: MAIDs (left part in each figure) and relevance graphs (right part) for one-shot simultaneous-move MARL interaction paradigms: (a) direct interaction, where agents interact directly; (b) global intervention, where a coordinator provides signals influencing all agents and their utilities; (c) targeted intervention, where a pre-strategy intervention applied on a targeted agent. In these diagrams, squares represent decision variables (specifically, red for the targeted agent, blue for other agent, yellow for a pre-decision); diamonds represent utility variables; and circles with $Z$ describe signals to guide agents for achieving their desired outcomes represented as utility variables. Relevance graphs depict strategic dependencies between these decision variables.
@@ 3.1 Formalizing MARL Interaction Paradigms
A MARL interaction paradigm can be specified as a MAID $(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U})$. $\mathcal{I}$ is the set of agents. $\mathcal{X}$ is the set of chance variables. $\mathcal{D}=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is the set of agent decision variables. $\mathcal{U}$ are utility variables representing objectives. The set of special chance variables ($\mathcal{Z}\subset\mathcal{X}$) describe signals to guide agents for achieving their desired outcomes.
Within the context of MARL, we define three interaction paradigms as follows:
(1) Direct Interaction. This paradigm represents the MARL where agents coordinate only via direct observation of the environment. There is no explicit, external coordination mechanism designed to steer agents towards their desired outcomes. Each agent must independently elicit a signal (a pattern of behaviours) based on local information to guide the other agent for achieving its desired outcome. For clarity of illustration, we focus on analyzing guidance to a single agent. We visualize this implicit process as a path from decision variables $D_{a}$ to the signal $Z$, and then to the utility nodes $U_{h}$, illustrated by MAID structures in Figure 1(a) and 2(a).
(2) Global Intervention. This paradigm features a centralized coordination providing simultaneous explicit guidance signals to all agents. In the MAID, as shown in Figure 1(b) and 2(b), this is demonstrated as information links from $Z$ to agents’ decision variables, and $Z$ usually influence their utility nodes (e.g., $U_{a}$ and $U_{h}$) directly, shaping their desired outcomes.
@@ 3.2 Interpreting MARL Interaction Paradigms
We now demonstrate how MAIDs interpret the three MARL interaction paradigms we mentioned above. Each MAID elucidates the structure of an interaction paradigm (through the role of variable $Z$) and its influence on decisions and utilities. Subsequently, we analyze how these structural differences translate into various strategic dependency patterns, as revealed by their relevance graphs.
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the direct interaction paradigm are cyclic, implying computational or theoretical intricacies in finding solutions (Remark 2.5). In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on targeted intervention provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention.
@@ 4.2 Solvability of MARL Learning Paradigms under Various Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms  [^3] Note that the IL algorithms we mention here exclude those with additional mechanisms designed to improve convergence, such as the optimistic mechanism (lauer2000algorithm, ). are the direct application of single-agent reinforcement learning algorithms, such as temporal-difference (TD) learning (sutton2018reinforcement, )[Chap. 6] and the actor-critic algorithms (konda1999actor, ). If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to independent learning, as discussed in Section 2.4. This aligns with the direct interaction paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^4] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, centralized training can be seen as an algorithmic approach to address the issues of pure IL algorithms existing in the direct interaction paradigm, rather than switching to other interaction paradigms.
@@ 5.1 Experimental Setups
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (direct interaction paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for independent learning and PQN-VDN for centralized training variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
@@ 6 Conclusion
Summary. We incorporate multi-agent influence diagrams (MAIDs) as a principled framework into designing and analyzing the targeted intervention paradigm in MARL. Our Pre-Strategy Intervention (PSI) as an implementation of the targeted intervention paradigm, applied to a single targeted agent, maximizes the causal effect on achieving a composite desired outcome that integrates both the primary task goal and an additional desired outcome. Furthermore, the relevance graphs of MAIDs offer theoretical insights into the solvability of MARL learning paradigms under various interaction paradigms. Experiments verify the effectiveness of the PSI.
@@ 3. Answer: [Yes]
  4. Justification: We raise our research question and solve it then evaluate.
@@ 13. Answer: [Yes]
  14. Justification: The omitted proof of main paper is in the Appendix.
@@ 18. Answer: [Yes]
  19. Justification: We provide implementation details in the Appendices.
@@ 28. Answer: [Yes]
  29. Justification: We provide details of the experimental settings in Appendices.
@@ 38. Answer: [Yes]
  39. Justification: We disclose the computational resources needed in Appendices.
@@ H.1 Implementation of Our Method
Target Agent Selection. For the experiments presented in this paper, we manually select a targeted agent in each environment. This experimental design allows for a clear and controlled analysis of the mechanism’s impact and mirrors practical scenarios where a human operator might manually choose a specific agent to intervene on.
@@ H.3 GNN Implementation
The foundation of our targeted intervention method (Pre-Strategy Intervention) is formed by the s-reachability graph criterion, introduced to identify the relevant policies of other agents for decision-making [16]. However, considering only policies is insufficient, as an agent’s policy may depend on other elements within the game [79]. Algorithm 1 provides an implementation how we can build connection with causal graph structure for pre-policy learning. By leveraging this graph structure, we incorporate prior knowledge about the game to help guide agents’ policy-making in practical. The feasibility of learning the causal graph during training has been demonstrated by Richens and Everitt [88], where agents can learn the causal model implicitly during interaction with the environment.

### v2

@@ Abstract
Steering cooperative multi-agent reinforcement learning (MARL) towards desired outcomes is challenging, particularly when the global guidance from a human on the whole multi-agent system is impractical in a large-scale MARL. On the other hand, designing external mechanisms (e.g., intrinsic rewards and human feedback) to coordinate agents mostly relies on empirical studies, lacking a easy-to-use research tool. In this work, we employ multi-agent influence diagrams (MAIDs) as a graphical framework to address the above issues. First, we introduce the concept of MARL interaction paradigms, using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we design a new MARL interaction paradigm, referred to as the targeted intervention paradigm that is applied to only a single targeted agent, so the problem of global guidance can be mitigated. In our implementation, we introduce a causal inference technique—referred to as Pre-Strategy Intervention (PSI)—to realize the targeted intervention paradigm. Since MAIDs can be regarded as a special class of causal diagrams, a composite desired outcome that integrates the primary task goal and an additional desired outcome can be achieved by maximizing the corresponding causal effect through the PSI. Moreover, the bundled relevance graph analysis of MAIDs provides a tool to identify whether an MARL learning paradigm is workable under the design of an MARL interaction paradigm. In experiments, we demonstrate the effectiveness of our proposed targeted intervention, and verify the result of relevance graph analysis.
@@ 1 Introduction
To address these challenges in MARL, researchers often introduce external mechanisms that guide agent interaction and learning—what we refer to as MARL interaction paradigms—to enhance performance, efficiency, and applicability. Intrinsic rewards, for instance, can improve sample efficiency and direct agent behaviour by encouraging exploration or coordination jaques2019social ; du2019liir ; mohamed2015variational . Another approach incorporating human feedback, leverages human expertise to accelerate learning. This helps agents tackle objectives that are complex to specify via simple reward functions, and aligns agent behaviour to desired norms or human values christiano2017deep ; hu2022human ; mnih2015human .
These MARL interaction paradigms cannot be fully captured by the conventional MARL learning paradigms (i.e., independent learning and centralized training).
@@
The main contributions of this paper are summarized as follows: (1) We introduce the concept of MARL interaction paradigms, using MAIDs to analyze and visualize both unguided self-organization and global guidance mechanisms in MARL. Then, we propose a new MARL interaction paradigm, referred to as the targeted intervention paradigm, where only a single targeted agent is intervened on. (2) The solvability of existing MARL learning paradigms under various MARL interaction paradigms can be analyzed, with the help of visualizing the corresponding relevance graphs. (3) Since MAIDs can be viewed as causal diagrams, we draw on the principle of causal inference to implement the targeted intervention paradigm, which we refer to as the Pre-Strategy Intervention (PSI). It is provable that the PSI can reach a composite desired outcome consisting of the primary task goal and an additional desired outcome. (4) The PSI is implemented as a pre-policy module that can be integrated into generic MARL algorithms.
@@ 3 Multi-Agent Reinforcement Learning Interaction Paradigms
We introduce the term MARL interaction paradigms to describe the ways in which agents’ interactions are structured—ranging from unguided self-organization to externally guided mechanisms such as intrinsic rewards or human feedback. Specifically, we first formalize two existing MARL interaction paradigms—self-organization and global intervention strouse2022collaborating ; mosqueira2023human ; wang2025m3hf —using multi-agent influence diagrams (MAIDs), and then introduce our proposed targeted intervention paradigm.
@@
Figure 1: MAIDs (left part in each figure) and relevance graphs (right part) for one-shot simultaneous-move MARL interaction paradigms: (a) self-organization, where agents coordinate with no external mechanism; (b) global intervention, where a coordinator provides signals influencing all agents and their utilities; (c) targeted intervention, where a pre-strategy intervention applied on a targeted agent. In these diagrams, squares represent decision variables (specifically, red for the targeted agent, blue for other agent, yellow for a pre-decision); diamonds represent utility variables; and circles with $Z$ describe signals to guide agents for achieving their desired outcomes represented as utility variables. Relevance graphs depict strategic dependencies between these decision variables.
@@ 3.1 Formalizing MARL Interaction Paradigms
An MARL interaction paradigm can be specified as an MAID $(\mathcal{I},\mathcal{X},\mathcal{D},\mathcal{U})$. $\mathcal{I}$ is the set of agents. $\mathcal{X}$ is the set of chance variables. $\mathcal{D}=\bigcup_{i\in\mathcal{I}}\mathcal{D}_{i}$ is the set of agent decision variables. $\mathcal{U}$ are utility variables representing objectives. The set of special chance variables ($\mathcal{Z}\subset\mathcal{X}$) describe signals to guide agents for achieving their desired outcomes.
We now define three MARL interaction paradigms as follows:
(1) Self-Organization. This paradigm refers to the MARL setting where agents coordinate solely via their direct observations of the environment, without any external guidance or structural intervention to steer them towards their desired outcomes. Each agent must independently elicit a signal (a pattern of behaviours) based on local information to guide the other agent for achieving its desired outcome. For clarity of illustration, we focus on analyzing guidance to a single agent. We visualize this implicit process as a path from decision variables $D_{a}$ to the signal $Z$, and then to the utility nodes $U_{h}$, illustrated by MAID structures in Figure 1(a) and 2(a).
(2) Global Intervention. This paradigm features a external, centralized coordination providing simultaneous explicit guidance signals to all agents. In the MAID, as shown in Figure 1(b) and 2(b), this is represented as information links from $Z$ to agents’ decision variables, and $Z$ usually influence their utility nodes (e.g., $U_{a}$ and $U_{h}$) directly, shaping their desired outcomes.
@@ 3.2 Interpreting MARL Interaction Paradigms
We now demonstrate how MAIDs interpret the three MARL interaction paradigms we mentioned above. Each MAID elucidates the structure of a paradigm (through the role of variable $Z$) and its influence on decisions and utilities. Subsequently, we analyze how these structural differences translate into various strategic dependency patterns, as revealed by their relevance graphs.
Analyzing the relevance graph (Definition 2.4) corresponding to each MAID, reveals the key solvability property [^2] Seeking equilibria according to processing decision variables in s-reachability order (influencers first) koller2003multi . of MARL learning algorithms. As shown in Figure 1(a) and 2(a), the relevance graphs for the self-organization paradigm are cyclic, implying computational or theoretical intricacies in finding solutions (Remark 2.5). In contrast, the relevance graphs for both the global intervention paradigm (Figures 1(b) and 2(b)) and targeted intervention paradigm (Figures 1(c) and 2(c)) are acyclic, suggesting they may be solvable by a broader class of MARL algorithms. It is possible in global intervention that agents learn independently based solely on the external central signal, in which case the relevance graph would have no edges at all, indicating no policy dependencies colas2022autotelicagentsintrinsicallymotivated . While both global intervention and targeted intervention paradigms offer acyclic structures, our proposed method relying on targeted intervention provides a distinct advantage: it achieves solvability and effectiveness only by intervention on a single targeted agent, offering a more practical approach in contrast to the global intervention.
@@ 4.2 Solvability of MARL Learning Paradigms under Various Interaction Paradigms
Independent Learning. Independent learning (IL) algorithms  [^3] Note that the IL algorithms we mention here exclude those with additional mechanisms designed to improve convergence, such as the optimistic mechanism (lauer2000algorithm, ). are the direct application of single-agent reinforcement learning algorithms, such as temporal-difference (TD) learning (sutton2018reinforcement, )[Chap. 6] and the actor-critic algorithms (konda1999actor, ). If a relevance graph is cyclic, it is not guaranteed to reach an NE by solely determining each agent’s decision variables with a generalized backward induction algorithm in analogy to independent learning, as discussed in Section 2.4. This aligns with the self-organization paradigm (Figure 3(a) and 4(a)), reflecting the non-stationarity dilemma (hernandez2019survey, ). Equipping IL algorithms with the global intervention (Figures 3(b) and 4(b)) and the targeted intervention (Figures 3(c) and 4(c)) paradigms can generate acyclic relevance graphs. Such acyclic relevance graphs can guarantee the tampered IL algorithms reaching an NE with theoretical guarantees.
Centralized Training. As discussed in Remark 2.5, the only way to address cyclic relevance graphs is enabling two agents’ decision making to be matched. One solution is transforming a cyclic relevance graph to a component graph, where each maximal strongly connected component (SCC) is regarded as a supernode de2023graph . More specifically, a maximal SCC consists of all the decision variables forming a cyclic relevance graph at each timestep. Koller and Milch koller2003multi showed that solving the acyclic component graph [^4] A component graph is always acyclic (cormen2022introduction, ). by the generalized backward induction algorithm can reach an NE. This is in coincidence associated with the centralized training employed to reach the maximum cumulative team rewards in Markov games (oliehoek2008optimal, ; littman2001value, ). In summary, centralized training can be seen as an algorithmic approach to address the issues of pure IL algorithms existing in the self-organization paradigm, rather than switching to other interaction paradigms.
@@ 5.1 Experimental Setups
Baselines. All baselines are implemented with the same architecture and training setups as our method. We evaluate our method against several base MARL algorithms referred to as Base MARL (self-organization paradigms): for MPE, these include DQN (mnih2015human, ), VDN (sunehag2017value, ) and QMIX (rashid2020monotonic, ); for Hanabi, we use IPPO (schulman2017proximal, ), MAPPO (yu2022surprising, ) and PQN (gallici2024simplifying, ) (with PQN-IQL for independent learning and PQN-VDN for centralized training variants). An ablation study that removes the pre-policy module while retaining the intrinsic reward maximization examines the effectiveness of the targeted intervention paradigm without the confounding effect of the pre-policy module. This ablated variant is referred to as Intrinsic Reward that still satisfies the targeted interaction paradigm, with removing pre-decision variables $D_{pre}$ in Figures 3 and 4 and connecting the $Z^{t}$ interpreted as the intrinsic rewards to maximize directly to the decision variables $D$. To emphasize the strength of our targeted intervention paradigm, we also compare against the existing approaches implemented in the global intervention paradigm, modifying LIIR du2019liir or LAIES liu2023lazy as representative policy-based or value-based methods. This baseline is referred to as Global Intervention.
@@ 6 Conclusion
Summary. We incorporate multi-agent influence diagrams (MAIDs) as a principled framework into designing and analyzing the targeted intervention paradigm in MARL. Our Pre-Strategy Intervention (PSI) approach as an implementation of the targeted intervention paradigm, applied to a single targeted agent, maximizes the causal effect on achieving a composite desired outcome that integrates both the primary task goal and an additional desired outcome. Furthermore, the relevance graphs of MAIDs offer theoretical insights into the solvability of MARL learning paradigms under various interaction paradigms. Experiments verify the effectiveness of the PSI.
@@ 3. Answer: [Yes]
  4. Justification: The key contributions of this paper are summarized in both the abstract and the final paragraph of the introduction, with each point corresponding to a dedicated section in the subsequent text.
@@ 13. Answer: [Yes]
  14. Justification: The omitted proof of main paper is in Appendix E.
@@ 18. Answer: [Yes]
  19. Justification: We provide implementation details in Appendix H.
@@ 28. Answer: [Yes]
  29. Justification: We provide details of the experimental settings in Appendices H and K.
@@ 38. Answer: [Yes]
  39. Justification: We disclose the computational resources needed in Appendix H.
@@ H.1 Implementation of Our Method
Targeted Agent Selection. For the experiments presented in this paper, we manually select a targeted agent in each environment. This experimental design allows for a clear and controlled analysis of the mechanism’s impact and mirrors practical scenarios where a human operator might manually choose a specific agent to intervene on.
@@ H.3 GNN Implementation
The foundation of our targeted intervention method (Pre-Strategy Intervention) is formed by the s-reachability graph criterion, introduced to identify the relevant policies of other agents for decision-making [16]. However, considering only policies is insufficient, as an agent’s policy may depend on other elements within the game [79]. Algorithm 1 provides an implementation how we can build connection with a causal graph structure for pre-policy learning. By leveraging this graph structure, we incorporate prior knowledge about the game to help guide agents’ policy-making in practical. The feasibility of learning the causal graph during training has been demonstrated by Richens and Everitt [88], where agents can learn the causal model implicitly during interaction with the environment.